After a completely inelastic collision, two objects of the same mass and same initial speed are found to move away together at 1/5 their initial speed. Find the angle between the initial velocities of the objects.

PLEASE WORK OUT. THANKS.
Answer #1

image from custom entry tool

let initial angle be θ1 and θ2

conservation og momentum along

mvcosθ1 + mvcosθ2 = 2mv/5

vcosθ1 + vcosθ2 = 2v/5

cosθ1 +cosθ2 = 2/5 --eqtn 1

vsinθ1 = vsinθ2

so θ1 = θ2

2cosθ1 = 2/5

cosθ1 =1/5

θ1 = 78.46 degree

so angle between their velocity = 2 * θ1 = 156.9 degree